83 less than a number equals -82
Answer by thepianist25(8)

From this word equation/expression, you can make an actual equation and solve for the unknown number. In this case, let's use x as our variable.
So, from the information provided:
x - 83 = -82
because,
"less than" means subtraction (-) and also means that you are taking the number 83 away from the variable, x;
and "is" means equals (=).
So, back to our equation:
x - 83 = -82
x - 83 + 83 = -82 + 83 (addition axiom)
x = 1
Checking the solution will show that the answer is correct.